An investor who has 1000 USD in hand and want to profit using the cross rate what is the profit?

SGD1.39/USD
THB 20.52/ SGD
THB 31.83/ USD
Answer #1

Start with 1000 USD
Convert to 31.83*1000 THB
Convert 31.83*1000 THB to 31.83*1000/20.52 SGD
Convert 31.83*1000/20.52 SGD to 31.83*1000/20.52*1/1.39 USD=1115.949346 USD

Net profit=1115.949346-1000=115.949346 USD